# Positive weak solutions of a double-phase variable exponent problem with a fractional-Hardy-type singular potential and superlinear nonlinearity

### v2

@@ 2 Mathematical Background and Preliminaries
$(N-\alpha q^{+})\int_{\Omega_{<}}\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx=(-q^{+})\int_{\Omega_{<}}|u|^{q^{+}-2}u\nabla u\cdot\frac{x}{|x|^{\alpha q^{+}}}dx.$ (7)
By Cauchy-Schwarz inequality, we have
$(N-\alpha q^{+})\int_{\Omega_{<}}\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx\leq q^{+}\int_{\Omega_{<}}\frac{|u|^{q^{+}-1}}{|x|^{\alpha q^{+}-1}}|\nabla u|dx.$ (8)
Then the Hölder inequality provides
$\int_{\Omega_{<}}\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx\leq\left(\frac{q^{+}}{N-\alpha q^{+}}\right)^{q^{+}}\int_{\Omega_{<}}|\nabla u|^{q^{+}}dx.$ (9)
By the continuous embedding $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow W_{0}^{1,q^{+}}(\Omega)$, we get
$\int_{\Omega_{<}}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq C_{N}(q^{+})\|u\|_{1,\mathcal{H},0}^{q^{+}}.$ (10)
For (b) and (c): $|x|^{\alpha}>|u(x)|$. Considering $(H_{1})$ again, for all $x\in\Omega_{<}$, $x\neq 0$, it reads
$\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\leq\frac{(1+\|\mu\|_{\infty})}{p^{-}}\frac{|u(x)|^{p^{-}}}{|x|^{\alpha p^{-}}}.$ (11)
Applying the same reasoning and steps of (a), it follows
$\int_{\Omega_{<}}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq C_{N}(p^{-})\|u\|_{1,\mathcal{H},0}^{p^{-}}.$ (12)
As a result of (10) and (12), it holds
$\int_{\Omega_{<}}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq C_{N}(p^{-},q^{+},\alpha)\|u\|_{1,\mathcal{H},0}^{\kappa},\,\,\ \forall u\in W_{0}^{1,\mathcal{H}}(\Omega_{<}),$ (13)
Case II: $x\in\Omega_{\geq}$.
In this case, we have $|x|\geq|x|^{\alpha}$, and hence we consider the following:
  • (a’)
    $|u(x)|\geq|x|\geq|x|^{\alpha}$,
  • (b’)
    $|x|>|u(x)|\geq|x|^{\alpha}$,
  • (c’)
    $|x|\geq|x|^{\alpha}>|u(x)|$.
Then for (a’) and (b’), we have $\frac{|u(x)|}{|x|^{\alpha}}\geq 1$; whereas, for (c’), we have $\frac{|u(x)|}{|x|^{\alpha}}<1$. Both cases follow in the same way as with Case I, so we can omit the details and write
$\int_{\Omega_{\geq}}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq C_{N}(p^{-},q^{+},\alpha)\|u\|_{1,\mathcal{H},0}^{\kappa},\,\,\ \forall u\in W_{0}^{1,\mathcal{H}}(\Omega_{\geq}),$ (14)
By combining (13) and (14), we can conclude that (1) holds for all $u\in W_{0}^{1,\mathcal{H}}(\Omega)$.
@@
$(ii)$ There exists $M>0$ such that $0<|x|<M$ for all $x\in\Omega\setminus\{0\}$. Thus, $\max\{|x|^{\alpha p(x)},|x|^{\alpha q(x)}\}<M^{\alpha\tau}$ which gives
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx>\frac{1}{M^{\alpha\tau}}\int_{\Omega}\left(|u|^{p(x)}+\mu(x)|u|^{q(x)}\right)dx>\frac{1}{M^{\alpha\tau}}\|u\|^{\kappa}_{\mathcal{H}}.$ (15)
In the sequel, we let $\mathcal{G}(u):=\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$.
Definition 2.10.
Let $X$ be a Banach space, and $J:X\rightarrow\mathbb{R}$ be a $C^{1}$-functional. We say that $J$ satisfies the Palais-Smale condition ($(\mathbf{PS})$ for short) if: every sequence $(u_{n})\subset X$ such that $J(u_{n})$ is bounded and $J^{\prime}(u_{n})\rightarrow 0$ in $X^{*}$ admits a convergent subsequence in $X$.
@@ 3 Variational Framework and Main Results
$\mathcal{I}(u)=\varrho_{\mathcal{H}}(u)+\mathcal{G}(u)-\lambda\int_{\Omega}F(x,u)dx.$
@@
$=\lim_{t\to 0}\int_{\Omega}\left(|x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\right$.
for all $u,\varphi\in W_{0}^{1,\mathcal{H}}(\Omega)$, $\gamma\in\mathbb{R}$. Using Young inequality, we obtain
$\bigg||x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\varphi\bigg|$ $\leq k|x|^{-\alpha p(x)}(|u|^{p(x)}+|\varphi|^{p(x)}),$ (19)
@@
$\bigg||x|^{-\alpha q(x)}\mu(x)|u+\varepsilon_{2}t\varphi|^{q(x)-2}(u+\varepsilon_{2}t\varphi)\varphi\bigg|$ $\leq k|x|^{-\alpha q(x)}(\mu(x)|u|^{q(x)}+\mu(x)|\varphi|^{q(x)}).$ (20)
where $k:=\frac{2^{q^{+}}(q^{+}-1)}{p^{-}}$. Using (19),(20) and Lemma 2.9, we can write
@@
$\leq k(|x|^{-\alpha p(x)}|u|^{p(x)}+|x|^{-\alpha q(x)}\mu(x)|\varphi|^{q(x)})+k(|x|^{-\alpha p(x)}|\varphi|^{p(x)}+|x|^{-\alpha q(x)}\mu(x)|\varphi|^{q(x)})$
$\leq kq^{+}C_{N}(p^{-},q^{+},\alpha)\left(\|u\|_{1,\mathcal{H},0}^{\kappa}+\|\varphi\|_{1,\mathcal{H},0}^{\kappa}\right).$ (21)
@@
Next, by Young inequality and Lemma 2.9, we obtain
@@
$\leq(q^{+}-1)\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\frac{|\varphi|^{p(x)}}{p(x)|x|^{\alpha p(x)}}\right)+\left(\frac{\mu(x)|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}+\frac{\mu(x)|\varphi|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$
$\leq(q^{+}-1)C_{N}(p^{-},q^{+},\alpha)\left(\|u\|_{1,\mathcal{H},0}^{\kappa}+\|\varphi\|_{1,\mathcal{H},0}^{\kappa}\right).$ (23)
@@
$\|\mathcal{G}^{\prime}(u)\|_{W_{0}^{1,\mathcal{H}}(\Omega)^{*}}=\sup_{\|\varphi\|_{1,\mathcal{H},0}\leq 1}|\langle\mathcal{G}^{\prime}(u),\varphi\rangle|\leq(q^{+}-1)C_{N}(p^{-},q^{+},\alpha)(\|u\|_{1,\mathcal{H},0}^{\kappa}+1),$ (24)
@@
$\left|\left|\xi\right|^{s-2}\xi-\left|\psi\right|^{s-2}\psi\right|\leq C_{s}\left|\xi-\psi\right|\{|\xi|+|\psi|\}^{s-2},\quad\xi,\psi\in\mathbb{R}^{N},\,\,s>1,$ (25)
@@
which implies that $\mathcal{I}(t\phi)\to$-$\infty$ as $t\to\infty$. Therefore, letting $\hat{u}=t\phi$, and hence $\|\hat{u}\|_{1,\mathcal{H},0}=t\|\phi\|_{1,\mathcal{H},0}$, leads to $\|\hat{u}\|_{1,\mathcal{H},0}>1>\gamma$ and $\mathcal{I}(\hat{u})<0$ provided $t$ is large enough,. ∎

### v3

@@ 2 Mathematical Background and Preliminaries
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx=-\frac{s}{N-\alpha s}\int_{\Omega}\mu(x)\,|u|^{s-2}u\nabla u\cdot\frac{x}{|x|^{\alpha s}}dx.$ (6)
Employing the Cauchy-Schwarz inequality gives
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx$ $\leq\frac{s}{N-\alpha s}\int_{\Omega}\mu(x)\,\frac{|u|^{s-1}}{|x|^{\alpha s-1}}|\nabla u|dx$
$=\frac{s}{N-\alpha s}\int_{\Omega}\mu(x)^{\frac{s-1}{s}}\,\frac{|u|^{s-1}}{|x|^{\alpha s-1}}\mu(x)^{\frac{1}{s}}\,|\nabla u|dx$ (7)
Applying Hölder’s inequality, and noticing that $(\alpha s-1)(\frac{s}{s-1})<\alpha s$, and assuming that $|x|<1$ (otherwise the result is trivial), gives
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx\leq$ $\frac{s}{N-\alpha s}\left(\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{(\alpha s-1)(\frac{s}{s-1})}}dx\right)^{\frac{s-1}{s}}\left(\int_{\Omega}\mu(x)\,|\nabla u|^{s}dx\right)^{\frac{1}{s}}$
$\leq\frac{s}{N-\alpha s}\left(\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx\right)^{\frac{s-1}{s}}\left(\int_{\Omega}\mu(x)\,|\nabla u|^{s}dx\right)^{\frac{1}{s}},$ (8)
which yields the following fractional-Hardy-type inequality
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx\leq H_{\mu,s}\int_{\Omega}|\nabla u|^{s}dx,$ (9)
where $H_{\mu,s}:=\|\mu\|_{\infty}\left(\frac{s}{N-\alpha s}\right)^{s}$ (we will let $H_{1,s}:=\left(\frac{s}{N-\alpha s}\right)^{s}$ if $\mu\equiv 1$). However, by the density argument, (9) holds for any $u\in W_{0}^{1,s}(\Omega)$.
Next, to ensure a rigorous analysis without overlapping conditions, we partition $\Omega$ into the following disjoint measurable sets:
$\Omega_{<}^{(a)}:=\Omega_{<}\cap\{|u(x)|\geq|x|^{\alpha}\},\quad\Omega_{<}^{(b)}:=\Omega_{<}\cap\{|u(x)|<|x|^{\alpha}\},\quad$
$\Omega_{\geq}^{(c)}:=\Omega_{\geq}\cap\{|u(x)|<|x|^{\alpha}\},\quad\Omega_{\geq}^{(d)}:=\Omega_{\geq}\cap\{|u(x)|\geq|x|^{\alpha}\},$ (10)
where
$\Omega_{<}:=\{x\in\Omega:|\nabla u(x)|<1\},\quad\Omega_{\geq}:=\{x\in\Omega:|\nabla u(x)|\geq 1\}.$ (11)
Since $\Omega$ is bounded and connected, we obviously have $\Omega=\Omega_{<}^{(a)}\cup\Omega_{<}^{(b)}\cup\Omega_{\geq}^{(c)}\cup\Omega_{\geq}^{(d)}$.
Step 1: Analysis on $\Omega_{<}$.
For $x\in\Omega_{<}$, using (9) we have
$\int_{\Omega_{<}}|\nabla u|dx\geq\int_{\Omega_{<}}|\nabla u|^{q^{+}}dx\geq\frac{1}{H_{\mu,q^{+}}}\int_{\Omega_{<}}\mu(x)\,\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx,$ (12)
and
$\int_{\Omega_{<}}|\nabla u|dx\geq\int_{\Omega_{<}}|\nabla u|^{p^{+}}dx\geq\frac{1}{H_{1,p^{+}}}\int_{\Omega_{<}}\frac{|u|^{p^{+}}}{|x|^{\alpha p^{+}}}dx.$ (13)
Estimate on $\Omega_{<}^{(a)}$: In this region we have $|u(x)|\geq|x|^{\alpha}$, and hence
$\int_{\Omega_{<}^{(a)}}\mu(x)\,\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx\geq\int_{\Omega_{<}^{(a)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx,$ (14)
and
$\int_{\Omega_{<}^{(a)}}\frac{|u|^{p^{+}}}{|x|^{\alpha p^{+}}}dx\geq\int_{\Omega_{<}^{(a)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx,$ (15)
Therefore,
$\int_{\Omega_{<}^{(a)}}|\nabla u|dx\geq\frac{1}{2H_{1}}\int_{\Omega_{<}^{(a)}}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx,$ (16)
where $\frac{1}{H_{1}}:=\frac{1}{H_{1,p^{+}}}\wedge\frac{1}{H_{\mu,q^{+}}}$.
Estimate on $\Omega_{<}^{(b)}$: In this region we have $|u(x)|<|x|^{\alpha}$, therefore
$\int_{\Omega_{<}}|\nabla u|dx\geq\int_{\Omega_{<}}|\nabla u|^{q^{-}}dx\geq\frac{1}{H_{\mu,q^{-}}}\int_{\Omega_{<}}\mu(x)\,\frac{|u|^{q^{-}}}{|x|^{\alpha q^{-}}}dx,$ (17)
and
$\int_{\Omega_{<}}|\nabla u|dx\geq\int_{\Omega_{<}}|\nabla u|^{p^{-}}dx\geq\frac{1}{H_{1,p^{-}}}\int_{\Omega_{<}}\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx.$ (18)
Since
$\int_{\Omega_{<}^{(b)}}\mu(x)\,\frac{|u|^{q^{-}}}{|x|^{\alpha q^{-}}}dx\geq\int_{\Omega_{<}^{(b)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx,$ (19)
and
$\int_{\Omega_{<}^{(b)}}\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx\geq\int_{\Omega_{<}^{(b)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx,$ (20)
we have
$\int_{\Omega_{<}^{(b)}}|\nabla u|dx\geq\frac{1}{H_{\mu,q^{-}}}\int_{\Omega_{<}^{(b)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx,$ (21)
and
$\int_{\Omega_{<}^{(b)}}|\nabla u|dx\geq\frac{1}{H_{1,p^{-}}}\int_{\Omega_{<}^{(b)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx.$ (22)
Therefore,
$\int_{\Omega_{<}^{(b)}}|\nabla u|dx\geq\frac{1}{2H_{2}}\int_{\Omega_{<}^{(b)}}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx,$ (23)
where $\frac{1}{H_{2}}:=\frac{1}{H_{1,p^{-}}}\wedge\frac{1}{H_{\mu,q^{-}}}$.
Step 2: Analysis on $\Omega_{\geq}$. For $x\in\Omega_{\geq}$, using (9) we have
$\int_{\Omega_{\geq}}|\nabla u|^{p(x)}dx\geq\int_{\Omega_{\geq}}|\nabla u|^{p^{-}}dx\geq\frac{1}{H_{\mu,p^{-}}}\int_{\Omega_{\geq}}\mu(x)\,\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx,$ (24)
and
$\int_{\Omega_{\geq}}|\nabla u|^{p(x)}dx\geq\int_{\Omega_{\geq}}|\nabla u|^{p^{-}}dx\geq\frac{1}{H_{1,p^{-}}}\int_{\Omega_{\geq}}\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx,$ (25)
Estimate on $\Omega_{\geq}^{(c)}$: In this region we have $|u(x)|<|x|^{\alpha}$. Since
$\int_{\Omega_{\geq}^{(c)}}\mu(x)\,\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx\geq\int_{\Omega_{\geq}^{(c)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx,$ (26)
and
$\int_{\Omega_{\geq}^{(c)}}\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx\geq\int_{\Omega_{\geq}^{(c)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx,$ (27)
it reads
$\int_{\Omega_{\geq}^{(c)}}|\nabla u|^{p(x)}dx\geq\frac{1}{H_{\mu,p^{-}}}\int_{\Omega_{\geq}^{(c)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx,$ (28)
and
$\int_{\Omega_{\geq}^{(c)}}|\nabla u|^{p(x)}dx\geq\frac{1}{H_{1,p^{-}}}\int_{\Omega_{\geq}^{(c)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx.$ (29)
Therefore,
$\int_{\Omega_{\geq}^{(c)}}|\nabla u|^{p(x)}dx\geq\frac{1}{2H_{3}}\int_{\Omega_{\geq}^{(c)}}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx,$ (30)
where $\frac{1}{H_{3}}:=\frac{1}{H_{1,p^{-}}}\wedge\frac{1}{H_{\mu,p^{-}}}$.
Estimate on $\Omega_{\geq}^{(d)}$: In this region, we have $|u(x)|\geq|x|^{\alpha}$. Considering the sufficiently large values of $|x|$, we have
$\int_{\Omega_{\geq}^{(d)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx\leq\|\mu\|_{\infty}\int_{\Omega_{\geq}^{(d)}}|u|^{q^{+}}dx,$ (31)
and
$\int_{\Omega_{\geq}^{(d)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx\leq\int_{\Omega_{\geq}^{(d)}}|u|^{q^{+}}dx.$ (32)
Hence,
$\int_{\Omega_{\geq}^{(d)}}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx\leq(1+\|\mu\|_{\infty})\int_{\Omega_{\geq}^{(d)}}|u|^{q^{+}}dx.$ (33)
Since $\Omega_{<}^{(a)},\Omega_{<}^{(b)},\Omega_{\geq}^{(c)}$ and $\Omega_{\geq}^{(d)}$ are measurable subsets of $\Omega$ with $\Omega=\Omega_{<}^{(a)}\cup\Omega_{<}^{(b)}\cup\Omega_{\geq}^{(c)}\cup\Omega_{\geq}^{(d)}$, we may extend all integrands by zero outside their respective sets in (16), (23), (30), and then integrate over all of $\Omega$, which gives
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx$
$\leq 2(H_{1}+H_{2})\int_{\Omega}|\nabla u|dx+2H_{3}\int_{\Omega}|\nabla u|^{p(x)}dx+(1+\|\mu\|_{\infty})\int_{\Omega}|u|^{q^{+}}dx.$ (34)
To conclude part $(i)$, we employ the embeddings in Proposition 2.7, which provides the desired result
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx\leq\hat{H}\|u\|_{1,\mathcal{H},0}^{\kappa},$
where
$\kappa=\begin{cases}1&\text{if }\,\|u\|_{1,\mathcal{H},0}<1,\\ q^{+}&\text{if }\,\|u\|_{1,\mathcal{H},0}\geq 1,\end{cases}$
and $\hat{H}:=2(H_{1}+H_{2})\hat{c}_{1}\vee 2H_{3}\hat{c}_{2}\vee(1+\|\mu\|_{\infty})\hat{c}_{3}$, where $\hat{c}_{1},\hat{c}_{2},\hat{c}_{3}$ are the corresponding embedding constants.
@@
$(ii)$ Since $\Omega$ is bounded, there exists $M>1$ such that $|x|<M$ for all $x\in\Omega$. Thus, for any $x\in\Omega\setminus\{0\}$, we have $|x|^{\alpha p(x)}<M^{\alpha p(x)}\leq M^{\alpha p^{+}}$ and $|x|^{\alpha q(x)}<M^{\alpha q(x)}\leq M^{\alpha q^{+}}$. Therefore, $|x|^{\alpha p(x)}\vee|x|^{\alpha q(x)}<M^{\alpha q^{+}}$, which gives
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx$ $>\frac{1}{M^{\alpha q^{+}}}\int_{\Omega}\left(|u|^{p(x)}+\mu(x)|u|^{q(x)}\right)dx$
$>\frac{1}{M^{\alpha q^{+}}}\|u\|^{{p^{-}\wedge q^{+}}}_{\mathcal{H}}.$ (35)
@@ 3 Variational Framework and Main Results
$\mathcal{I}(u)=\varrho_{\mathcal{H}}(u)+\mathcal{G}(u)-\lambda\int_{\Omega}F(x,u)dx,$
where $\mathcal{G}(u):=\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$.
@@
$=\lim_{t\to 0}\int_{\Omega}\left(|x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\varphi\right$.
$+\left.|x|^{-\alpha q(x)}\mu(x)|u+\varepsilon_{2}t\varphi|^{q(x)-2}(u+\varepsilon_{2}t\varphi)\varphi\right)dx,$
for all $u,\varphi\in W_{0}^{1,\mathcal{H}}(\Omega)$, $\gamma\in\mathbb{R}$. Using the Young inequality, we obtain
$\bigg||x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\varphi\bigg|$ $\leq k_{1}|x|^{-\alpha p(x)}(|u|^{p(x)}+|\varphi|^{p(x)}),$ (39)
@@
$\bigg||x|^{-\alpha q(x)}\mu(x)|u+\varepsilon_{2}t\varphi|^{q(x)-2}(u+\varepsilon_{2}t\varphi)\varphi\bigg|$ $\leq k_{2}|x|^{-\alpha q(x)}(\mu(x)|u|^{q(x)}+\mu(x)|\varphi|^{q(x)}).$ (40)
where $k_{1}:=\frac{2^{p(x)-1}(p(x)-1)+1}{p(x)}$, and $k_{2}:=\frac{2^{q(x)-1}(q(x)-1)+1}{q(x)}$.
Using (39),(40) and Lemma 2.10, we can write
@@
$\leq k_{1}(|x|^{-\alpha p(x)}|u|^{p(x)}+|x|^{-\alpha q(x)}\mu(x)|u|^{q(x)})+k_{2}(|x|^{-\alpha p(x)}|\varphi|^{p(x)}+|x|^{-\alpha q(x)}\mu(x)|\varphi|^{q(x)})$
$\leq k\hat{H}\left(\|u\|_{1,\mathcal{H},0}^{\kappa}+\|\varphi\|_{1,\mathcal{H},0}^{\kappa}\right),$ (41)
where $k:=\frac{2^{q^{+}}(q^{+}-1)}{p^{-}}\vee k_{1}\vee k_{2}$.
@@
Next, by the Young inequality and Lemma 2.10, we obtain
@@
$\leq\frac{q^{+}}{p^{-}}\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\frac{|\varphi|^{p(x)}}{p(x)|x|^{\alpha p(x)}}\right)+\left(\frac{\mu(x)|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}+\frac{\mu(x)|\varphi|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$
$\leq\frac{q^{+}}{p^{-}}\hat{H}\left(\|u\|_{1,\mathcal{H},0}^{\kappa}+\|\varphi\|_{1,\mathcal{H},0}^{\kappa}\right).$ (43)
@@
$\|\mathcal{G}^{\prime}(u)\|_{W_{0}^{1,\mathcal{H}}(\Omega)^{*}}$ $=\sup_{\|\varphi\|_{1,\mathcal{H},0}\leq 1}|\langle\mathcal{G}^{\prime}(u),\varphi\rangle|\leq\frac{q^{+}}{p^{-}}\hat{H}\left(\|u\|_{1,\mathcal{H},0}^{\kappa}+1\right),$ (44)
@@
$\left|\left|\xi\right|^{m-2}\xi-\left|\psi\right|^{m-2}\psi\right|\leq C_{m}\left|\xi-\psi\right|\{|\xi|+|\psi|\}^{m-2},\quad\xi,\psi\in\mathbb{R}^{N},\,\,m>1,$ (45)
@@
Note that $(\mathbf{AR})$ stands for the Ambrosetti–Rabinowitz condition [2].
@@
which implies that $\mathcal{I}(t\phi)\to$-$\infty$ as $t\to\infty$. Therefore, letting $\hat{u}=t\phi$, and hence $\|\hat{u}\|_{1,\mathcal{H},0}=t\|\phi\|_{1,\mathcal{H},0}$, leads to $\|\hat{u}\|_{1,\mathcal{H},0}>1>\gamma$ and $\mathcal{I}(\hat{u})<0$ provided $t$ is large enough. ∎
Definition 3.5.
Let $X$ be a Banach space, and $J:X\rightarrow\mathbb{R}$ be a $C^{1}$-functional. We say that $J$ satisfies the Palais-Smale condition ($(\mathbf{PS})$ for short) [2] if: every sequence $(u_{n})\subset X$ such that $J(u_{n})$ is bounded and $J^{\prime}(u_{n})\rightarrow 0$ in $X^{*}$ admits a convergent subsequence in $X$.